[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Mar 17, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Thu Mar 17 13:10:11 UTC 2016


Here is today's summary of NHC computer operations:
--- Monica

----------

TAFB:
None
HSU:
None
Other items:
1. Disabling of USB drives will continue for the next few days. Operational  
windows systems will be completed by EOB today, how ever non-operational  
systems may require everyone to work one on one with Kim Gordon.

2. NCO is still working to replace the unisys server to restore access to  
the radar data

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160317/5d5cf5cb/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list